OpenStudy (anonymous):

DEFG Is a rectangle. DF=5x-5 and EG=x+11. find the value of x and the length of each diagonal. (Draw and label the sketch below)

OpenStudy (anonymous):

df=eg x=4

OpenStudy (anonymous):

DF and EG are diagonals. Diagonals in a rectangle are always equal to each other, so we can set 5x-5=x+11 add 5 to both sides 5x=x+16 subtract x from both sides 4x=16 divide both sides by 4 x=4

OpenStudy (anonymous):

what does DF and Eg equal/?

OpenStudy (anonymous):

15